# A Spatio-Temporal Machine Learning Model for Mortgage Credit Risk: Default Probabilities and Loan Portfolios

### v1

@@ 4 Conclusion
We introduce a novel machine learning model combining tree-boosting with a latent spatio-temporal Gaussian process that accounts for frailty correlation. We compare our proposed model to an independent linear hazard model, a linear model with spatial frailty effects, and a linear model with spatio-temporal frailty effects and find that predictive default probabilities and predictive loan portfolio loss distributions obtained with our novel model are more accurate. Using interpretability tools for machine learning models, we provide evidence that the reasons for this better performance are interactions and non-linear effects in the predictor variables that cannot be captured by a linear functional form. In addition, we find that there are relatively strong spatio-temporal frailty effects.
@@ Acknowledgments
We would like to thank Raffaella Calabrese for her helpful advice for gathering the mortgage credit risk data. This research was partially supported by the Swiss Innovation Agency - Innosuisse (grant number ‘55463.1 IP-ICT’).
@@ Appendix A Appendix
A.1 Additional results for prediction of individual default probabilities
Figure 10: Temporal out-of-sample test H-measure (higher = better).
A.2 Additional results for prediction of loan portfolio loss distributions
A.3 Posterior mean for the latent Gaussian process of the linear spatio-temporal model
A.4 Tuning parameters
Tuning parameter Candidate values
Number of trees {1,2,…,1000}
Learning rate {10,1,0.1}
Maximal tree depth {2,3,5,10}
Minimal number of samples per leaf {10,100,1000}
L2 regularization {0,1,10}
Table 4: Candidate tuning parameters for the tree-boosted spatio-temporal frailty model.
Figure 14: Selected tuning parameters.
A.5 Summary statistics of predictor variables
@@
orig_dti 1.000 27.000 34.679 34.339 41.000 65.000
orig_cltv 2.000 66.000 80.000 75.735 88.000 534.000
orig_upb 9000.000 106000.000 160000.000 186612.656 243000.000 1581000.000
cnt_ltv 0.000 59.252 72.373 69.203 79.351 526.587

### v2

@@ 4 Conclusion
We introduce a novel machine learning model combining tree-boosting with a latent spatio-temporal Gaussian process that accounts for frailty correlation. We compare our proposed model to an independent linear hazard model, a linear model with spatial frailty effects, and a linear model with spatio-temporal frailty effects and find that predictive default probabilities and predictive loan portfolio loss distributions obtained with our novel model are more accurate. Although the differences in prediction accuracy for individual default probabilities are relatively small, they are nonetheless important, as even small differences are relevant for large portfolios of mortgages. For instance, during the global financial crisis around the year 2008, we observe large differences in predictive loan portfolio loss distributions with mean losses of frailty machine learning models being almost US \$1 billion more accurate compared to linear models. Using interpretability tools for machine learning models, we find that there are strong interactions and non-linear effects in the predictor variables. This is further evidence that linear models likely assume a misspecified functional form. In addition, we find that there are sizable spatio-temporal frailty effects.
A potential limitation of the spatio-temporal models we considered in this article is that they overestimate the portfolio loss in the year 2021 after the COVID-19 pandemic which lead to a rather peculiar one-year peak in default probabilities during the year 2020. On the other hand, linear and machine learning models that only account for spatial frailty do not have this drawback. Future research can investigate whether and how such an overestimation can be mitigated, for instance, using other forms of spatio-temporal models. In general, future research can systematically analyze and compare different space-time covariance functions for spatio-temporal frailty modeling.
For confidentiality reasons, the exact locations of the mortgaged properties are not included in the single-family loan-level data set from Freddie Mac we used in this article. Consequently, spatial and spatio-temporal frailty correlation could only be modeled at a relatively coarse resolution. Having exact coordinates could potentially further improve the accuracy of spatial and spatio-temporal frailty models. Future research is required to investigate whether modeling of spatial and spatio-temporal frailty correlations using more granular credit risk data sets leads to more accurate predictions. We expect potential improvements not just for mortgages but in general when modeling loans associated with a specific spatial location such as SME, credit card, student, and peer-to-peer loans.
@@ Acknowledgments
We would like to thank Raffaella Calabrese for her helpful advice for gathering the mortgage data. We also thank Pascal Gantenbein and Christian Kleiber for their useful feedback. This research was partially supported by the Swiss Innovation Agency - Innosuisse (grant number ‘55463.1 IP-ICT’).
@@ Appendix A Appendix
A.1 Vecchia approximations for the latent Gaussian process
We apply a Vecchia approximation to the distribution of the latent Gaussian variable $b\in\mathbb{R}^{R}$:
$\begin{split}p(b|\theta)&=\prod_{r=1}^{R}p(b_{r}|(b_{1},\dots,b_{r-1}),\theta)\\ &\approx\prod_{r=1}^{R}p(b_{r}|b_{N(r)},\theta),\end{split}$
where $b_{N(r)}$ are subsets of the conditioning sets $(b_{1},\dots,b_{r-1})$, and $N(r)$ denotes the corresponding subsets of indices. If $r>\tilde{m}+1$, $N(r)$ is often chosen such that it contains the indices of the $\tilde{m}$ nearest neighbors of $b_{r}$ in the conditioning set. Otherwise, $N(r)$ equals $(1,\dots,r-1)$. Using the properties for conditional multivariate normal distributions, we get:
$p(b_{r}|b_{N(r)},\theta)=\mathcal{N}(A_{r}b_{N(r)},D_{r}),$
where
$\begin{split}A_{r}&=\Sigma_{r,N(r)}\Sigma_{N(r)}^{-1}\\ D_{r}&=\Sigma_{r,r}-\Sigma_{r,N(r)}\Sigma_{N(r)}^{-1}\Sigma_{N(i),i}.\end{split}$
$\Sigma\in\mathbb{R}^{R\times R}$ is the covariance matrix of $b$ and $\Sigma_{r,N(r)}\in\mathbb{R}^{1\times\tilde{m}}$ denotes the row vector that contains the entries of the $r$-th row of $\Sigma$ that have a column index in $N(r)$, and $\Sigma_{N(r)}\in\mathbb{R}^{\tilde{m}\times\tilde{m}}$ denotes the matrix that contains the entries of $\Sigma$ that have a row and column index in $N(r)$. Defining the lower triangular matrix $B\in\mathbb{R}^{R\times R}$ with $1$’s on the diagonal, off-diagonal entries
$B_{r,N_{(r)}}=-A_{r},$
and 00 otherwise, and the diagonal matrix $D\in\mathbb{R}^{R\times R}$ with $D_{i}$ on the diagonal, we obtain an approximate distribution for the latent Gaussian variable:
[…]
The corresponding precision matrix is sparse and given by $\tilde{\Sigma}^{-1}=B^{T}D^{-1}B$. Calculating the Cholesky factor of $\Sigma_{N(r)}$ is the main computational burden of the approximation. Thus, a Vecchia approximation costs $\mathcal{O}(R\times\tilde{m}^{3})$, and saving $B$ requires $O(R\times\tilde{m})$ memory storage. Often, accurate approximations are obtained for small $\tilde{m}$’s.
A.2 Annual spatial default rates
Figure 10: Annual spatial default rates. No default rates are shown for the grey areas because there are too few loan-year observations.
A.3 Summary statistics of predictor variables
@@
orig_dept_to_income 1.000 27.000 34.679 34.339 41.000 65.000
orig_loan_to_value 2.000 66.000 80.000 75.735 88.000 534.000
orig_unpaid_balance 9000.000 106000.000 160000.000 186612.656 243000.000 1581000.000
curr_loan_to_value 0.000 59.252 72.373 69.203 79.351 526.587
@@
gdp_growth -15.019 2.539 3.994 3.951 5.578 25.184
income_per_capita 9.937 10.522 10.693 10.699 10.867 11.490
expend_per_capita 9.723 10.332 10.484 10.477 10.621 11.282
unemployment_rate 2.100 4.400 5.400 6.021 7.300 13.500
hpi_growth -23.070 -0.600 3.460 3.099 6.450 31.710
construction_growth -33.149 -0.570 5.345 3.724 8.895 42.042
op_surplus_growth -22.903 2.150 4.630 4.561 6.569 34.548
inflation_rate -0.356 1.465 2.069 2.147 3.157 4.698
djia_growth -33.840 -0.610 7.260 7.449 18.820 26.500
@@
A.4 Tuning parameters
Tuning parameter Candidate values
Number of trees {1,2,…,1000}
Learning rate {10,1,0.1}
Maximal tree depth {2,3,5,10}
Minimal number of samples per leaf {10,100,1000}
L2 regularization {0,1,10}
Table 6: Candidate tuning parameters for the tree-boosted spatial and spatio-temporal frailty models.
Figure 11: Selected tuning parameters.
A.5 Runtimes
Time (s)
Linear independent 15.03
Linear spatial 11.54
Linear spatio-temporal 224.48
LaGaBoost spatial 49.34
LaGaBoost spatio-temporal 1637.72
Table 7: Average wall clock times (s) for training on data up to the year 2014.
A.6 Additional results for prediction of individual default probabilities
Figure 12: Temporal out-of-sample test H-measure (higher = better).
Compared models P-value
LaGaBoost spatio-temporal vs. LaGaBoost spatial 1.27e-02
LaGaBoost spatio-temporal vs. Linear spatio-temporal 7.56e-27
LaGaBoost spatio-temporal vs. Linear spatial 1.83e-07
LaGaBoost spatio-temporal vs. Linear independent 6.76e-09
LaGaBoost spatial vs. Linear spatio-temporal 7.78e-09
LaGaBoost spatial vs. Linear spatial 2.94e-03
LaGaBoost spatial vs. Linear independent 1.24e-03
Linear spatio-temporal vs. Linear spatial 2.24e-09
Linear spatio-temporal vs. Linear independent 1.63e-08
Linear spatial vs. Linear independent 3.05e-01
Table 8: P-values of the pairwise DeLong tests using the function roc.test in the R package pROC. AUCs are based on the pooled predictive probabilities of each model over all prediction years.
A.7 Additional results for prediction of loan portfolio loss distributions
A.8 Posterior mean for the latent Gaussian process of the linear spatio-temporal model